OpenStudy (anonymous):

What are the solutions for x in the proportion 5x-6/x=2x/2

OpenStudy (anonymous):

5x-6/x=2x/2 5x^2-6=x^2 5x^2-x^2=6 4x^2=6 x^2=6/4=3/2 \[x=\pm \sqrt{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

x = −2 and x = −3 x = 2 and x = 3

OpenStudy (anonymous):

x =3

OpenStudy (anonymous):

whitch one of those would that amount to?

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

i think is x=3